OpenStudy (anonymous):

36x^2+25y^2-z^2-60xy

zepdrix (zepdrix):

\[\large 36x^2-60xy+25y^2-z^2 \qquad=\qquad (6x)^2-2(6x)(5y)+(5y)^2-z^2\] Hmm looks like we can simplify down the x and y's! :)

zepdrix (zepdrix):

\[\large \color{royalblue}{a^2-2ab+b^2=(a-b)^2}\]

OpenStudy (anonymous):

the problem is what I should do with the z^2

zepdrix (zepdrix):

leave the z alone :D

OpenStudy (anonymous):

Thank you

OpenStudy (unklerhaukus):

the z's are all together already , they cant be simplified further